Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 54 points. They won 12 more games than they tied. How many wins and how many ties did the team have?
HOw many wins did the team have?
How many ties did the team have?

Answer by checkley77(12569)

(12+X)2+X=54
24+2X+X=54
3X-54-24
3X=30
X=30/3
X=10 GAMES TIED.
10+12=22 GAMES WON.
PROOF:
2*22+10*1=54
44+10=54
54=54